You just turned 23 years old. your aunt will give you $ 1000 for your birthday. you will use the money to open a bank account. How much do you expect to have in the account on your 65th birthday?

you just turned 23. your aunt will give you $ 1000 for your birthday. You will use the money to open a bank account. The account has an annual return of 12%. How much do you expect to have in the account on your birthday?

you just turned 23. your aunt will give you $ 1000 for your birthday. You will use the money to open a bank account. The account has an annual return of 12%. How much do you expect to have in the account on your 65 birthday?
Answer #1

So the aunt pays $1000 at the age of 23 with which a savings account is opened that gives 12% annual return.

The amount expected in the account on the 65th birthday could be calculated using the Compound Interest formula

A = P * (1 + r)n

Where:

  • A = Accrued Amount on 65th birthday (principal + interest)
  • P = Principal Amount = $1000
  • r = Annual Nominal Interest Rate as a decimal = 0.12
  • n = 65 -23 = 42

therefore A = 1000 * (1+ 0.12)^42 = $ 116,723.14

116723.137